• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Saha

    Log entry time 16:02:34 on November17,2009

    Entry number 300576

    keyword=Shift Summary

    Shift Summary

    0800: Production Running LD2 Target 90 uA


    Run: 14015/25555/5103 100k events
    Run: 14016/25556/5104 100k events

    0950: Start Moller measurements


    30: After moving Hall A target to EMPTY position MCC has problems with Hall A target vacuum level.

    1125: MCC ask for Controlled Access to check Hall A target vacuum valve.

    1145: Beam permit in Hall A

    1210: LHRS Dipole and RHRS Q1 and Q3 going down. Controlled access. Jack going to hall to reset PS.

    1230: Hall in Beam permit. Proceeding with Moller measurement.

    1410: Hall in Controlled access. Heidi to reset LHRS Q3 and Sasha to reset IOCHLA for Moller target.

    1600: Getting to do a Moller measurement